OpenStudy (anonymous):

HELPPP PLEASE !!!!! Show all steps to solved the problem Be sure you check your solution for correctness 5n+6=-4

OpenStudy (anonymous):

GOOD ONE

OpenStudy (smarty):

5n+6=-4 Subtract 6 from both sides 5n=-4-6 5n=-10 Divide both sides by 5 n=-2 5(-2)+6=-4 -10+6=-4 -4=-4

OpenStudy (kaederfds):

5n=-4-6 5n=-10 n=-2

OpenStudy (anonymous):

Thanx (: